OpenStudy (anonymous):

Solve: 2x\equiv 3(mod 5)

jimthompson5910 (jim_thompson5910):

Can you solve 2x = 1 (mod 5) ?

OpenStudy (anonymous):

nooo lol

jimthompson5910 (jim_thompson5910):

does 2*1 = 1 (mod 5)?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

what about 2*2 = 1 (mod 5)?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

and what about 2*3 = 1 (mod 5)?

OpenStudy (anonymous):

ok x=3

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

So 2*3 = 1 (mod 5) We want that right side to be a 3 instead of a 1, so multiply both sides by 3

OpenStudy (anonymous):

y?

jimthompson5910 (jim_thompson5910):

2*3 = 1 (mod 5) 3*2*3 = 3*1 (mod 5) 2*(3*3) = 3 (mod 5) 2*9 = 3 (mod 5) 2*4 = 3 (mod 5)

jimthompson5910 (jim_thompson5910):

because at the top you wanted to solve 2x = 3 (mod 5)

OpenStudy (anonymous):

ohhh i get it

jimthompson5910 (jim_thompson5910):

yes, you start with 2x = 1 (mod 5), then you scale appropriately

OpenStudy (anonymous):

2*9 = 3 (mod 5) 2*4 = 3 (mod 5) how did u do this?

jimthompson5910 (jim_thompson5910):

9 = 4 (mod 5)

jimthompson5910 (jim_thompson5910):

since they both have the same remainder after dividing by 5

jimthompson5910 (jim_thompson5910):

9/5 = 1 remainder 4 4/5 = 0 remainder 4

OpenStudy (anonymous):

okkkk gotcha

jimthompson5910 (jim_thompson5910):

that's great

OpenStudy (anonymous):

so x=4 not 3

jimthompson5910 (jim_thompson5910):

yes because 2*4 = 8 = 3 (mod 5)

OpenStudy (anonymous):

okkkkkk thank youuuuuuuu

jimthompson5910 (jim_thompson5910):

you're welcome